Write the slope-intercept equation for the line with the given slope that contains the given point.
m=8/3; (3,-4)

Answer by stanbon(75887)

Write the slope-intercept equation for the line with the given slope that contains the given point.
m=8/3; (3,-4)
------
Form: y = mx +b
Solve for "b" when m = 8/3, x = 3, and y = -4
---
-4 = (8/3)*3 + b
-4 = 8 + b
b = -12
------
Equation:
y = (8/3)x - 12
====================
Cheers,
Stan H.
==========